OpenStudy (anonymous):

x+5y less than or equal to -10

OpenStudy (anonymous):

\[ x + 5y \le -10 \]And the question is ... ?

OpenStudy (anonymous):

graph it

OpenStudy (anonymous):

can u solve for y?

OpenStudy (anonymous):

Isolate the y-term and divide: \[ 5y \le -10 - x \]\[ \boxed {y \le -2 - \dfrac {x}{5}}. \]Then, graph the line as if it were = and shade below.